40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 5 15:28:38 2014, Steve, Update, SUS, HeNe laser test preparation beforeRSMqpdMoved.jpgafterSRMqpdMoved.jpg1103Ptest.pngND1onQPD.png
    Reply  Fri Sep 5 16:25:36 2014, Steve, Update, SUS, center ITMY oplev 
    Reply  Tue Sep 9 08:22:28 2014, Steve, Update, SUS, HeNe laser test preparation HeNecompared.png
       Reply  Tue Sep 9 16:59:32 2014, Steve, Update, SUS, HeNe laser test preparation 1103PqpdTranss.png
          Reply  Wed Sep 10 14:58:58 2014, Jenne, Update, SUS, OpLev test: New channels 
             Reply  Thu Sep 11 17:12:42 2014, Steve, Update, SUS, OpLev test: old SP qpd connected 
                Reply  Fri Sep 12 16:30:48 2014, Steve, Update, SUS, HeNe laser test preparation ND2.pngqpd222ND21103P.jpg
                   Reply  Wed Sep 17 13:41:00 2014, Jenne, Update, SUS, New calibrated channels for test QPD NewQPDcalibratedChans.png
                      Reply  Thu Sep 18 09:17:47 2014, Steve, Update, SUS, HeNe laser test  calibrated_16hrs.pngQPD222cal.jpg1103P40s40min.png
                         Reply  Mon Sep 22 10:14:07 2014, Steve, Update, SUS, HeNe laser test  1103Pmon3d.png
                            Reply  Tue Sep 23 08:34:14 2014, Steve, Update, SUS, HeNe laser test  tempDrop.png
                               Reply  Thu Sep 25 10:53:32 2014, Steve, Update, SUS, HeNe laser test  centeredDay6.png
          Reply  Wed Sep 10 15:31:16 2014, Steve, Update, SUS, HeNe laser test preparation SRMqpdisBack0nSRMsus.pngslowVSfast.png
    Reply  Wed Oct 15 17:09:49 2014, manasa, Update, General, Diode laser test preparation 
       Reply  Wed Oct 29 18:07:28 2014, manasa, Update, General, Diode laser test preparation 
Message ID: 10513     Entry time: Wed Sep 17 13:41:00 2014     In reply to: 10504     Reply to this: 10517
Author: Jenne 
Type: Update 
Category: SUS 
Subject: New calibrated channels for test QPD 

Steve asked about calibrating the QPD, so I set up some new epics records so that we can have calibrated versions of the QPD output.

The new channels are called C1:ASC-TESTQPD_Y_Calc and C1:ASC-TESTQPD_X_Calc for pitch and yaw, respectively.


Details:

 * I modified /cvs/cds/caltech/target/c1iscaux/QPD.db to add 2 new channels.  Since we are currently plugged into the IPPOS channels, I didn't want to modify the units of IPPOS, which is why I created new channels.  The new channels are just the IPPOS normalized X and Y channels, multiplied by a calibration factor.  Steve has already done a rough calibration for his setup, so I used those numbers (0.15 urad/ct for pitch and 0.25 urad/ct for yaw). 

* Rebooted c1iscaux.  This required adding it to chiara's /etc/hosts file.

* Added the channels to the /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini file so that the channels would appear in dataviewer.

* Restarted the framebuilder daqd process.


How to modify the calibration:

1) On a control room workstation, cd /cvs/cds/caltech/target/c1iscaux to get to the right folder.  (Note that this is still in the old cvs/cds place, *not* the new opt/rtcds place)

2)  open the epics database file by typing sudo emacs QPD.db.  Since this is a protected file, you need to use the "sudo" command, and will have to type in the usual controls password. 

3)  Find the "records" that have the channel names C1:ASC-TESTQPD_Y_Calc and C1:ASC-TESTQPD_X_Calc by scrolling down.  (Right now they are on lines #550 and #561 of the text file).

4) For each of these 2 records, modify the calibration in the line that says something like field(CALC,"(A*0.25)").  In this example, the current calibration is 0.25 urad/oldCount.  Change the number to the new value.

5) Save the file.  If you followed the procedure in step2 and used the emacs program and you can't use the mouse, do the following:  Hold down the "ctrl" key.  Keeping ctrl pushed down, push the "x" key.  Still keeping ctrl pushed down, push the "s" key. 

6) Close the file.  If you followed the procedure in step2 and used the emacs program and you can't use the mouse, do the following:  Hold down the "ctrl" key.  Keeping ctrl pushed down, push the "x" key.  Still keeping ctrl pushed down, push the "c" key. 

7) Reboot the slow computer called c1iscaux.  You should be able to do this remotely by typing telnet c1iscaux, and then typing reboot.  If that doesn't work, you may have to go into the IFO room and power cycle the crate by turning the key.  This computer is in 1Y3, near the bottom.

8) Check that you can see your channels - you should be finished now!

For steps 3 and 4, here is a screenshot of the lines in the text file:

NewQPDcalibratedChans.png

ELOG V3.1.3-